# Two Types of AI Existential Risk: Decisive and Accumulative

### v2

@@ 6 The perfect storm MISTER
Consider the highly interconnected world of 2035 where the pervasive integration of AI and IoT technologies has transformed almost every aspect of daily life. [^16] The Internet of Things (IoT) refers to the network of physical devices embedded with sensors, software, and other technologies for the purpose of connecting and exchanging data with other devices and systems over the Internet. These devices range from ordinary household items like refrigerators and thermostats to sophisticated industrial tools. Internet of things represents the idea of a highly interconnected world where real-time data exchange and automation are pervasive. Cities embody a higher level of automation as compared to today, various functions in domestic and industrial sectors are automated, and even the most mundane devices like mirrors and refrigerators have become part of a vast, data-exchanging network. Frontier AI systems have become the backbone of this connected world. However, beneath the surface of this technological connectivity, vulnerabilities and risks have been brewing.
@@
AI-driven manipulation. The capability of AI to create convincing deepfakes reaches a critical height, where distinguishing fact from deepfake-generated reality becomes nearly impossible. [^17] Concerns regarding AI-generated fake realities, especially in the context of manipulation and misinformation, have long been topics of research. However, recent advancements in generative AI are introducing new dimensions to this issue. As shown in various sources, including a recent post by Chase Dean on Twitter, there is an increasing possibility of using deepfake technology to craft highly convincing but completely fabricated representations of public figures, events, or news stories. Currently, no methods are entirely reliable in distinguishing these fabrications from actual reality. This capability opens the door to the creation of hyper-personalized, AI-generated propaganda and similar persuasive narratives, which can be strategically leveraged for social engineering purposes — manipulating public opinion, influencing electoral outcomes, inciting civil unrest, diminishing oppositional viewpoints, and exploiting existing social and epistemic divides.
@@
Proliferation of Insecurity threats. In the domestic sphere, everyday personal IoT devices like smart mirrors and refrigerators transcend their roles as mere conveniences. They now represent significant points of vulnerability. With frontier AI capabilities, cybercriminals can penetrate these devices in a more expansive way, leading to widespread identity theft and ushering in a new era of digital espionage. What were initially perceived as isolated incidents of breaches gradually coalesce into a discernible pattern, signifying a more amplified erosion of digital security.
@@
The expansion of IoT devices also paves the way for the creation of extensive, interconnected botnets. These AI-powered networks, once relatively benign, show agentic abilities and become capable of launching unprecedented Distributed Denial of Service attacks, against critical infrastructures, including national power grids and communication networks. They represent a significant leap in cybersecurity risks. Each attack, incrementally more sophisticated than the last, signifies a disturbing escalation from individual cybersecurity concerns to widespread threats against national security.
@@
Moreover, as AI technologies becomes more widely available, they facilitate the emergence of new forms of bioterrorism. Private research labs with minimal expertise in synthetic biology and chemistry are now using AI to develop more infectious and deadly pathogens. The dual-use nature of AI in biotechnology — its potential for both beneficial and harmful applications — initially envisioned for medical breakthroughs, is maliciously repurposed to engineer biological weapons. [^18] The integration of AI and drug discovery, especially in the context of developing toxic substances or biological agents, is already a significant wake-up call. A notable example of this call is the empirical research conducted by Collaborations Pharmaceuticals, Inc., which investigated the feasibility of creating harmful biochemical agents based on VX-like compounds (Urbina et al., 2022). This research highlights how the integration of machine learning models with specialized knowledge in fields like chemistry or toxicology can substantially lower technical barriers in generation of bio-weapons. Tools like retrosynthesis software, which assist in the design of molecules by reversing their synthetic processes, exemplify this trend.
Global mass surveillance and erosion of trust. By 2035, the implications of AI-induced mass surveillance have profoundly transformed socio-political structures worldwide. The earlier instances, such as China’s social credit system and the controversial NSA’s mass data collection, were just the precursors to a now-global trend. Governments across various political spectra have embraced advanced AI for mass surveillance, frequently under the guise of national security. This landscape is marked by an array of pervasive surveillance systems, some overt and others clandestinely operated under national security pretexts.
In this future context, the reach of government surveillance have gradually eroded the sanctity of privacy even in putative democracies. The constant monitoring and evaluation of citizens by AI systems instill a culture of self-censorship among citizens, diluting the diversity of thought and social trust crucial for healthy democracy. In several countries, this leads to a radical erosion of trust in media as well as public and private institutions.
Economic destabilization. In the financial market infrastructure, AI algorithms increasingly play a substantial role in analyzing and simulating market trends. This progression, however, introduces new challenges to economic stability. One emerging concern is the use of AI in creating sophisticated, but deceptive financial products. These products, while not outright phantom instruments, are complex derivatives or investment vehicles. It becomes increasingly difficult to distinguish these products from more traditional, stable investments, leading to potential market disruptions. The impact of these AI-enhanced financial products is particularly significant in already vulnerable economies facing systemic issues and cyber threats. The introduction of these complex investment tools further undermines trust in financial systems, intensifying the risk of economic downturns.
Additionally, the potential for AI to unintentionally trigger market instability becomes a realistic concern. Automated trading algorithms, responding rapidly to market signals, now can inunprecedented speed amplify market volatility at unprecedented speeds. This phenomenon is exemplified in incidents like flash crashes — i.e., sudden and drastic market fluctuations driven by the high-speed responses of automated trading systems.
@@
Rights infringement. The pervasive application of AI in mass surveillance and extensive data collection practices is increasingly encroaching upon basic human rights. Privacy breaches have become alarmingly routine as AI systems gather and analyze personal data on an unprecedented scale. This constant monitoring undermines the right to privacy, a cornerstone of individual freedom. Furthermore, these surveillance mechanisms exert a chilling effect on freedom of expression. Individuals, aware of the omnipresent AI-driven surveillance, may self-censor or refrain from expressing dissenting opinions, leading to a stifling of public discourse and democratic engagement. The situation is compounded by AI systems that enable discriminatory profiling and unwarranted scrutiny of individuals. Such practices, often lacking transparency and accountability, lead to systemic unjust treatment and exacerbate existing societal inequalities at scale.
@@
In the perfect storm MISTER Scenario, a series of interconnected AI-induced risks coalesce into a catastrophic sequence of events, each exacerbating the next, leading to a potential existential crisis for humanity.
@@
In this scenario, the x-catastrophe emerges from the synergistic failure of systems critical to the functioning and survival of human civilization. The simultaneous and compounded nature of these crises creates a perfect storm situation where not only is recovery extremely challenging, but the potential for irreversible collapse is a stark reality.
Before discussing the ramifications of the accumulative AI x-risk hypothesis, let’s first consider and address two potential objections.
@@ 7 Objections and replies
Objection 1: The societal breakdown or accumulative civilizational collapse, as it is characterized, does not equate to the x-catastrophes often envisioned in AI risk discussions, such as an ASI power-seeking entity deliberately aims to destruct humanity. Historically, civilizational collapses have happened and, while significant, have not equated to human extinction or unrecoverable collapse.
Reply 1: This objection, while historically grounded, may not fully consider the unique AI x-catastrophic threats posed in our modern, interconnected global civilization. The modern world functions as a tightly interconnected global system, far more integrated than any past civilization. The COVID-19 pandemic serves as a contemporary example, where a health crisis originating in one region quickly escalated into a global emergency, disrupting economies, supply chains, and everyday life worldwide. The perfect storm MISTER example illustrates how a collapse in one sector or region can have rapid, global impacts in today’s interconnected world, a phenomenon not seen in past civilizational collapses.
@@
Reply 2: The complexity inherent in the accumulative AI x-risk model is actually a key advantage. It promotes a comprehensive understanding of the diverse impacts of AI across societal, economic, and ecological domains. By acknowledging the interdependencies and potential for cascading effects within these systems, we can increase resilience and adaptability against x-catastrophic disruptions, leading to the development of more robust long-term strategies.
@@
While the paperclip maximizer example offers a simple and elegant model, it deeply overlooks the nuanced realities of AI’s interactions with complex global systems. The accumulative model, in contrast, acknowledges this complexity and provides a more realistic portrayal of potential risks. It emphasizes detailed, empirically-grounded analysis, which is crucial for identifying critical areas sensitive to existential risks.
@@
Future efforts should aim to further validate and refine the accumulative hypothesis through system dynamic simulations (e.g., Karnopp et al. (2012)). These simulations would offer tangible insights into the complex interplay of AI-induced risks, enhancing our understanding and ability to model these interactions accurately. Although this paper does not resolve all the questions surrounding the accumulative model, it shows the need for ongoing research and deeper exploration in the effective management of AI x-risks. The approach towards a more nuanced understanding of AI’s potential impact on our world is both necessary and urgent.
8 Implications
In this section, I briefly discuss five governance implications of this expanded conception of AI x-risks.
8.1 Consolidated approach to AI risk governance
Recognizing the distinct nature of decisive and accumulative AI x-risks calls for a holistic evaluation of how we govern AI risks.
The governance of decisive ASI x-risks may be best served by centralized and stringent control measures. These controls aim to quickly address and mitigate the immediate and widespread consequences that such particular risks could unleash. For instance, stringent regulation of ASI research and development is needed. This might include international agreements akin to the Nuclear Non-Proliferation Treaty, ensuring that advancements towards ASI are closely monitored and controlled. Rapid-response protocols, similar to those used in handling biochemical threats, could be implemented.
In the event of an ASI-related crisis, a centralized authority could quickly mobilize resources and expertise to mitigate the threat. On the other hand, accumulative AI x-risks, which develop gradually and are often less perceptible in the short term, calls for decentralized, adaptable governance. This could involve a network of local and regional AI oversight bodies that monitor the impacts of AI in their respective areas, akin to environmental protection agencies that oversee local environmental concerns.
The unification of different domains in AI risk governance not only streamlines efforts but also avoids the unnecessary costs associated with rebranding. AI development is a gradual and continuous process, and similarly, the development of tools for detecting and mitigating AI risks should also be progressive and sustained. The integrated approach prevents the redundant effort of ’reinventing the wheel’ for each distinct aspect of AI x-risk mitigation.
To clarify, the consolidated approach to AI risk governance does not necessitate a direct link between every project and each facet of social risks to x-risks. Rather, it advocates for a meta-view or a macro-level perspective. This perspective facilitates a comprehensive understanding of how various projects, while not necessarily directly interlinked, collectively contribute to the overarching objective of mitigating AI x-risks. In light of this discussion, I propose to establish a dynamic ”risk contribution checklist.” This checklist, designed to be both flexible and regularly updated, would serve as a tool for researchers and institutions engaged in AI development. It would enable a coordinated approach to risk mitigation, ensuring that diverse projects align with and contribute effectively to the broader goal of managing AI existential risks.
8.2 Harmonization of ethical and existential risks

### v3

@@ 4 The perfect storm MISTER
Consider the highly interconnected world of 2040 where the pervasive integration of AI tools, AI assistants, AI agents, and Internet of Things (IoT) technologies has transformed almost every aspect of daily life. [^24] IoT refers to the network of physical devices embedded with sensors, software, and other technologies for the purpose of connecting and exchanging data with other devices and systems over the Internet (Rose et al., 2015; Li et al., 2015). These devices range from ordinary household items like refrigerators and thermostats to sophisticated industrial tools. Internet of things represents the idea of a highly interconnected world where real-time data exchange and automation are pervasive. Cities embody a higher level of automation as compared to today, with sustainability assistants (Rillig and Kasirzadeh, 2024) optimizing resource usage, AI agents (Xi et al., 2023) managing various functions in domestic and industrial sectors, and even the most mundane devices like mirrors and refrigerators have become part of a vast, data-exchanging network. Personalized AI assistants (Gabriel et al., 2024) have become the backbone of this connected world, serving roles from decision-making algorithms to social companions. However, beneath the surface of this technological connectivity, vulnerabilities and risks have been brewing.
@@
Manipulation by AI assistants and agents. The abuse and misuse of AI systems for creating convincing deepfakes and misinformation reaches a critical point, where the information ecosystem becomes so polluted that rational public discourse becomes nearly impossible. [^25] Concerns regarding AI-generated fake realities, especially in the context of manipulation and misinformation, have long been topics of research (Whittaker et al., 2020; Sharma and Kaur, 2022). However, recent advancements in generative AI are introducing new dimensions to this issue. As shown in various sources, including a recent post by Chase Dean on Twitter, there is an increasing possibility of using deepfake technology to craft highly convincing but completely fabricated representations of public figures, events, or news stories. Currently, no methods are entirely reliable in distinguishing these fabrications from actual reality. The manipulation architecture operates through “cognitive cascade captures” (Hazrati and Ricci, 2024; Deldjoo et al., 2024) where initial successful manipulation creates vulnerabilities for subsequent influence attempts.
Advanced AI technologies have enabled the creation of hyper-personalized propaganda and persuasive narratives, which can be strategically leveraged for social engineering purposes such as manipulating group identities, amplifying existing prejudices, exploiting belief systems through synthetic evidence, and creating personalized epistemic bubbles.
Unlike traditional static influence attempts, AI assistants and agents embedded in recommender systems could maintain consistent manipulation pressure while dynamically adapting to individual and collective response patterns (Kasirzadeh and Evans, 2023; Carroll et al., 2024). This unprecedented technological capability, in turn, poses fundamental challenges for maintaining individual and collective autonomy, preserving shared reality frameworks, protecting democratic discourse, and ensuring societal resilience.
Insecurity threats. The proliferation of IoT devices in domestic environments has fundamentally transformed the security landscape, creating unprecedented vulnerabilities in personal digital spaces.
@@
Digital-security threats. Smart devices, from mirrors to refrigerators, have evolved beyond their roles as mere conveniences to significant points of vulnerability. Cybercriminals can now penetrate these devices in increasingly sophisticated ways, leading to widespread identity theft and ushering in a new era of digital espionage. What were initially perceived as isolated breaches have gradually coalesced into a discernible pattern, signifying a more profound erosion of digital security.
@@
The expansion of IoT devices has simultaneously paved the way for the creation of extensive, interconnected botnets. These AI-powered networks, once relatively benign, now demonstrate agentic abilities and have become capable of launching unprecedented Distributed Denial of Service attacks against critical infrastructures, including national power grids and communication networks. Each attack, incrementally more sophisticated than the last, represents a disturbing escalation from individual cybersecurity concerns to widespread threats against national security.
@@
Bio-security threats. As AI technologies become more widely available, they facilitate the emergence of new forms of bioterrorism. Private research labs with minimal expertise in synthetic biology and chemistry are now using AI to develop more infectious and deadly pathogens. The dual-use nature of AI in biotechnology — its potential for both beneficial and harmful applications — initially envisioned for medical breakthroughs, is maliciously repurposed to engineer biological weapons. [^26] The integration of AI and drug discovery, especially in the context of developing toxic substances or biological agents, is already a significant wake-up call. A notable example of this call is the empirical research conducted by Collaborations Pharmaceuticals, Inc., which investigated the feasibility of creating harmful biochemical agents based on VX-like compounds (Urbina et al., 2022). This research highlights how the integration of machine learning models with specialized knowledge in fields like chemistry or toxicology can substantially lower technical barriers in generation of bio-weapons. Tools like retrosynthesis software, which assist in the design of molecules by reversing their synthetic processes, exemplify this trend.
Epistemic insecurity. Advanced AI assistants and agents have introduced fundamental challenges to both public and private epistemic infrastructures. [^27] Following Milano and Prunkl (2024) and Kay et al. (2024), epistemic infrastructure includes the systems, institutions, and practices through which knowledge claims are verified, transmitted, and maintained within societies. In public domains, this has manifested through the systematic erosion of shared verification mechanisms, where traditional epistemic authorities face unprecedented challenges as synthetic content becomes increasingly indistinguishable from authentic documentation. This disruption strikes at the heart of societal knowledge validation systems, undermining established protocols for information verification and authentication.
In the private sphere, personal communication channels—traditionally resistant to large-scale manipulation through social trust mechanisms—have become increasingly vulnerable to synthetic content that convincingly mimics trusted sources (Pennycook and Rand, 2021). This disruption extends beyond mere communication interference, affecting personal correspondence authenticity, private record verification, and the fundamental reliability of interpersonal trust mechanisms. The impact on private epistemic frameworks represents a significant shift in how individuals verify and validate information within their personal networks.
Surveillance and erosion of Trust. The transformation of mass surveillance through AI represents one of the deepest shifts in the relationship between state power and individual liberty. What began as isolated initiatives have evolved into a global phenomenon that transcends traditional political classifications, fundamentally altering the perceptions of privacy, social cohesion, and democratic governance.
The historical trajectory of mass surveillance has reached a troubling convergence between authoritarian and democratic governance models. Early warning signs emerged with China’s social credit system and the NSA’s PRISM program, but these now appear almost wide-spread. [^28] For an analysis of this evolution, see Lyon (2021) examination of how pandemic-era surveillance measures accelerated the adoption of AI-driven monitoring systems. Multiple earlier revelations have particularly warned against the erosion of democratic norms by AI-enabled surveillancce technologies such as spywares (Farrow, 2022; Rujevic, 2024). [^29] Farrow’s groundbreaking 2022 investigation exposed how putative democracies have embraced surveillance technologies traditionally associated with authoritarian regimes. The cases of Greece’s phone-hacking campaign targeting opposition politicians and Poland’s deployment of Pegasus spyware against civil society actors demonstrate how surveillance technologies can be weaponized even within democratic frameworks. The European Parliament’s 2024 Special Committee report on surveillance spyware documents numerous instances of democratic governments using these technologies against their own citizens.
The societal implications of ubiquitous surveillance manifest in what Han (2015) describes as the “transparency society,” where the mere possibility of observation fundamentally alters social behavior. [^30] Han’s (2015) The Transparency Society provides a philosophical framework for understanding how constant surveillance reshapes social relations and individual psychology. Earlier studies (Kaminski and Witnov, 2014) investigated and document situations where citizens modify their behavior not in response to actual surveillance but to its perceived omnipresence. [^31] Murray et al.’s (2024) empirical study documents self-censorship across multiple societies, correlating directly with the implementation of AI surveillance systems. This chilling effect on public behavior and discourse represents a particularly insidious threat to democratic vitality, as it operates through self-imposed constraints rather than direct coercion.
Economic destabilization. By 2040, the global economy has entered an unprecedented phase of instability, driven by the rapid and unmanaged deployment of AI systems across industries since the late 2020s. Nearly 40% of pre-2025 jobs have been eliminated by automation, creating massive structural unemployment. The pace and sophistication of AI adoption left no sufficient time for meaningful workforce transitions. The promised creation of new jobs never materialized at scale — AI systems became in charge of handling their own maintenance, optimization, and even creative development. The proposed solution of Universal Basic Income (UBI), while theoretically promising, has fallen short in practice due to political constraints and corporate resistance to the taxation necessary for meaningful implementation. [^32] See Auken (2016). The concentration of AI capabilities among a handful of tech conglomerates has exacerbated wealth inequality to historic levels, with the top 0.1% now controlling over 70% of global wealth.
The economic upheaval has been amplified by the fragmentation of the global order into competing digital-industrial blocs. The US-led Western alliance and the China-centered Asian sphere have created parallel technological and financial ecosystems, effectively splitting the world economy. This digital iron curtain has disrupted decades of global trade integration, with companies forced to maintain separate systems and standards for each bloc. The emergence of competing digital currencies has undermined the dollar-based financial system, while AI-powered economic warfare — including automated sanctions enforcement, algorithmic trade restrictions, and digital blockades — has become a daily reality. Smaller nations find themselves forced to align with one bloc or risk economic isolation, further destabilizing regional powers and trade relationships.
Market stability faces additional challenges from the acceleration of algorithmic trading systems. While high-frequency trading is not new, the integration of advanced AI capabilities introduces novel forms of systemic risk. These systems can now process and react to market signals at unprecedented speeds, potentially creating feedback loops that amplify market volatility. The phenomenon extends beyond simple flash crashes to what might be termed ”cascade failures,” where AI-driven trading systems interact in ways that create emergent instability patterns.
@@
Rights infringement. The pervasive application of AI in mass surveillance and extensive data collection practices has extensively encroached upon basic human rights. Privacy breaches have become alarmingly routine as AI systems gather and analyze personal data on an unprecedented scale. This constant monitoring undermines the right to privacy, a cornerstone of individual freedom. Furthermore, these surveillance mechanisms exert a chilling effect on freedom of expression. Individuals, aware of the omnipresent AI-driven surveillance, may self-censor or refrain from expressing dissenting opinions, leading to a stifling of public discourse and democratic engagement. The situation is compounded by AI systems that enable discriminatory profiling and unwarranted scrutiny of individuals. Such practices, often lacking transparency and accountability, lead to systemic unjust treatment and exacerbate existing societal inequalities at scale.
@@
In the perfect storm MISTER Scenario, against this background condition, a series of interconnected AI-induced risks coalesce into a catastrophic sequence of events, each exacerbating the next, leading to an existential crisis for humanity.
@@
In this scenario, the x-catastrophe arise from the synergistic failure of systems critical to the functioning and survival of human civilization. The simultaneous and compounded nature of these crises creates a perfect storm situation where not only is recovery extremely challenging, but the potential for irreversible collapse is a stark reality.
With the perfect storm MISTER scenario established, we now compare the causal pathways underlying decisive and accumulative hypotheses from a systems analysis perspective.
5 Pathway to decisive ASI x-risk
The decisive ASI x-risk hypothesis examines scenarios, like the paperclip maximizer, where ASI could abruptly trigger existential catastrophes. This hypothesis assumes catastrophic outcomes arise from a single, identifiable cause: the creation of a misaligned superintelligence capable of rapid system-wide disruption. First, ASI serves as the initial perturbation source, introducing novel disruptions into the global system. Second, the extensive connectivity of modern global infrastructures to ASI enables rapid propagation of ASI-initiated disruptions. Third, predominantly unidirectional dependencies prevent the system from self-correcting, instead reinforcing and accelerating the catastrophic trajectory.
Figure 2: Pathway to decisive ASI x-catastrophe
Figure 2 illustrates the causal pathway in decisive AI x-risk scenarios. Assumptions about the type of pathway for risk propagation are as follows. Modern civilization operates through densely connected subsystems — from financial markets and supply chains to communication infrastructures and social institutions. In the figure, square nodes represent various subsystems in the global world, with red squares indicating subsystems impacted by the ASI and grey squares showing those not yet impacted. Stronger impacts are shown with bold red links, while weaker or potential connections are indicated in grey. The arrangement follows a temporal sequence for illustrative purposes.
In scenarios like the paperclip maximizer, what begins as a simple optimization goal (maximize paperclips) cascades through multiple subsystems as the ASI pursues instrumental subgoals like resource acquisition. Initial control of computational resources spreads through economic networks, infrastructure control cascades through technological systems, and responses to human resistance ripple through social and political networks. Unlike natural systems that have inherent balancing mechanisms (like predator populations limited by prey scarcity), an ASI creates self-reinforcing cycles without effective counterbalances. For example, the ASI’s acquisition of computing resources, as an instrumental subgoal, increases its capabilities, enabling more effective strategies for pursuing further subgoals. Each cycle amplifies the ASI’s ability to pursue its ultimate objective while simultaneously reducing the possibility of external intervention. With access to critically important components (from global manufacturing to energy grids), the ASI can leverage all necessary subsystems, with no significant subsystem immune from its influence (illustrated by the red diamond with x). This progression lacks natural checks - nothing inherently constrains its escalating influence as it pursues its final goal of converting resources into paperclips.
6 Pathway to accumulative AI x-risk
The accumulative AI x-risk hypothesis posits that existential risks may arise from multiple interacting disruptions that compound over time, progressively weakening systemic resilience until a triggering event causes unrecoverable collapse. Unlike decisive scenarios with a single causal path, this hypothesis involves multiple causal processes that collectively contribute to existential risk.
First, different types of AI systems serve as initial perturbation sources, where localized impacts, even if minor at the outset, can aggregate and intensify across various subsystems. Second, while modern infrastructures are highly interconnected, AI systems typically impact specific subsystem clusters — creating selective pathways for disruptions to propagate through the network rather than the pervasive reach by ASI seen in decisive AI x-risk scenarios. Third, disruptions from different AI clusters interact and amplify each other as they spread through connected subsystems, creating cumulative effects larger than their initial impacts. Finally, as impacts from multiple AI clusters accumulate across subsystems, the system’s capacity for self-correction diminishes, making each new disruption more likely to reinforce rather than resolve existing instabilities.
Figure 3: Pathway to accumulative AI x-catastrophe
Figure 3 illustrates the causal pathway in accumulative AI x-risk scenarios. Each hexagon represents a cluster of problems compounded by AI assistants or agents in perfect storm MISTER: Manipulation by AI assistants and agents, Insecurity threats, Surveillance and erosion of Trust, Economic destabilization, and Rights infringement. Within the hexagons, circles and squares represent different entities (institutions, humans, communities) interacting with AI systems.
The accumulative AI x-risk manifests itself from the potential of these interlinked and reciprocal changes to progressively destabilize key subsystems. Although no single AI application poses an immediate existential threat, the aggregated impact of AI across various adapted subsystems leads to critical systemic imbalances or crises (diamonds in red). [^33] One major challenge is to develop dynamic models that capture these shifting feedback mechanisms and to validate these models through both historical data and controlled experimentation. There remains uncertainty about how different types of feedback loops interact and how interventions can be designed to maintain a system within safe operational boundaries. Open questions pertain to the identification of early warning signals that might indicate a system is approaching a critical threshold where negative feedback can no longer contain an escalating process. Next steps include applying system dynamics to study the proposed concepts more concretely.
To summarize, the decisive and accumulative AI x-risk hypotheses differ in three key aspects. First, in perturbation source: a single ASI versus multiple AI systems creating localized disruptions. Second, in propagation pattern: the ASI’s pervasive reach enabling rapid system-wide cascades versus selective pathways through specific subsystem clusters. Third, in catastrophic development: immediate unidirectional acceleration versus gradual accumulation of interacting disruptions that progressively degrade system resilience. While both pathways can lead to system-wide catastrophe, they do so through fundamentally different causal mechanisms.
Before examining the implications of the accumulative AI x-risk hypothesis for governance and long-term safety, let us address several potential objections.
@@ 7 Objections and replies
Objection 1: The societal breakdown or accumulative civilizational collapse, as characterized here, does not equate to the x-catastrophes often envisioned in AI risk discussions, such as an ASI power-seeking entity deliberately aiming to destroy humanity. Historically, civilizational collapses have occurred and, although significant, have not been equated to human extinction or unrecoverable collapse.
Reply 1: This objection, while historically grounded, may not fully consider the unique AI x-catastrophic threats posed in our modern, interconnected global civilization. The modern world functions as a tightly interconnected global system, far more integrated than any past civilization. The COVID-19 pandemic serves as a contemporary example, where a health crisis originating in one region quickly escalated into a global emergency, disrupting economies, supply chains, and everyday life around the world. The perfect storm MISTER example illustrates how a collapse in one sector or region can have rapid, global impacts in today’s interconnected world, a phenomenon not seen in past civilizational collapses.
@@
Reply 2: The complexity inherent in the accumulative AI x-risk model is actually a key advantage. By acknowledging the interdependencies and potential for cascading effects across societal, economic, and ecological domains, we can develop more targeted monitoring mechanisms. These mechanisms would track critical thresholds, measure compounding effects, and identify early warning signals of system destabilization. While such monitoring systems require significant development, they are essential for understanding how AI impacts accumulate across complex global systems.
@@
While the paperclip maximizer example offers a simple and elegant model, it deeply overlooks the nuanced realities of AI’s interactions with complex global systems. The accumulative model, in contrast, acknowledges this complexity and provides a more realistic portrayal of potential risks. It calls for detailed, empirically-grounded monitoring and analysis, which is crucial for identifying critical areas sensitive to existential risks.
@@
In addition to establishing monitoring mechanisms, future efforts should aim to further validate and refine the accumulative hypothesis through system dynamic simulations (e.g., Karnopp et al. (2012)). These simulations would offer tangible insights into the complex interplay of AI-induced risks. Although this paper does not resolve all the questions surrounding the accumulative model, it shows the need for ongoing research and deeper exploration in the effective conceptualization and management of AI x-risks.
8 Risk governance and accumulative risk hypothesis
The imperative to govern AI has become almost synonymous with mitigating its risks (Kaminski, 2023). The risk-based approach to AI governance is evident in major policy frameworks such as the NIST AI Risk Management Framework (National Institute of Standards and Technology, 2024), which presents risk taxonomies that align with earlier classifications (Bender, 2013; Weidinger et al., 2021; Bird et al., 2023) and regulatory approaches like the EU AI Act’s risk-based categorization.
But how do different conceptions of AI risk fit together in risk governance efforts? The risk landscape is fragmented across multiple taxonomies with calls for governing social and ethical risks, catastrophic risks (Kasirzadeh, 2024), extreme risks (Shevlane et al., 2023; Bengio et al., 2024), and existential risks. These separate terminologies create what we might term “the risk fragmentation problem” — where distinct approaches to conceptualizing AI risk fail to provide complete coverage, leaving dangerous blind spots where risks can accumulate unnoticed.
This fragmentation manifests in several ways. Social and ethical risk frameworks operate in isolation from frameworks addressing catastrophic risks (Anthropic, 2023). Immediate risk assessments rarely connect with analyses of long-term societal implications. Assessment methodologies remain siloed within their specific domains and terminologies, creating spaces where risks go unmonitored or unaddressed. [^36] There are at least four principal models for risk governance, each representing different balances between quantitative rigor and democratic accountability (Kaminski, 2023). First, the U.S. administrative model (1960s-1980s) emphasizes heavily quantitative risk assessment methodologies, prioritizing cost-benefit analysis and measurable outcomes (Boyd, 2012). This approach, while offering analytical precision, often struggles to account for uncertainties and non-quantifiable or hardly-quantifiable risks that characterize emerging new technologies. [^34] This tension between quantification and uncertainty becomes particularly acute when dealing with novel AI risks that lack historical precedent for statistical analysis. Second, the democratic oversight model, exemplified by the National Environmental Policy Act (NEPA), emphasizes public participation and transparent decision-making processes (Froomkin, 2015; Kaminski and Malgieri, 2020). This approach can incorporate precautionary principles, acknowledging that when facing potentially catastrophic risks, the absence of complete scientific certainty should not preclude protective measures. [^35] The precautionary approach becomes especially relevant for AI governance given the potential for irreversible societal impacts. Third, the centralized risk evaluator assesses risk on a macro-level (Hampton, 2005; Black and Baldwin, 2010). Regulators identify the risk to be managed, select a level of risk tolerance, assess the harms and the likelihood of their occurrence, assign risk scores to firms or activities (such as “high,” “medium,” or “low”), and link the allocation of enforcement and inspection resources to risk scores. The Draft EU AI Act is a clear descendant of this kind of law. Fourth, enterprise risk management approaches, as represented by NIST standards (National Institute of Standards and Technology, 2024) and frontier model safety frameworks (Kasirzadeh, 2024), focus on how companies can organize internally to mitigate their risks. They may conduct ongoing risk analysis and mitigation to avoid liability or other penalties, whether regulatory or market-based. While enterprise risk management can occur in the absence or shadow of law, regulators can also participate by nudging companies to conduct risk mitigation through oversight, through the threat of regulatory enforcement, by offering safe harbors, or by issuing best practices or other guidance. Enterprise risk management is typically (1) cyclical and ongoing, and (2) organizational in nature. These models diverge not only in their approaches to hard versus soft law but also in their mechanisms for ensuring accountability and their treatment of uncertainty. While quantitative cost-benefit analysis often dominates traditional risk assessment, the unique challenges posed by AI technologies suggest the need for hybrid approaches that can incorporate both precautionary principles and rigorous analytical methods. Recent developments in AI risk assessment suggest a growing recognition of the need to combine multiple governance approaches to address the full spectrum of potential risks.
The accumulative conception of AI x-risk — or its weaker interpretation as accumulative AI catastrophic risk — provides a unifying framework that can help bridge fragmented risk governance approaches. By emphasizing how risks compound and interact across domains, this perspective incentivizes connecting previously siloed risk frameworks.
8.1 Holistic approach to AI x-risk governance
The distinction between decisive and accumulative AI x-risks requires different, complementary, governance approaches.
Decisive ASI x-risks call for centralized control measures similar to nuclear non-proliferation frameworks. This includes international monitoring of advanced AI development, strict development protocols, and coordinated emergency response mechanisms. The potential for rapid, system-wide impacts necessitates unified oversight and quick response capabilities.
Accumulative x-risks, by contrast, require distributed monitoring systems that can track how multiple AI impacts compound across different domains. This suggests a network of oversight bodies monitoring specific sectors and subsystems, while sharing data about emerging risk patterns. Like financial regulators tracking systemic risk, these bodies would need mechanisms to detect when accumulated disruptions approach critical thresholds.
These different governance needs can be integrated through a tiered framework: distributed monitoring for accumulative risks, coupled with centralized oversight for advanced AI development. This approach leverages existing governance structures while adding new capabilities for tracking risk accumulation.
8.2 Unifying social and existential risks